We take a 10-year mortgage for $300,000 at 7.25% p.a. It is to be repaid in monthly repayments.

a. What is the repayment amount? Assume that interest is compounded monthly. Which formula should you use to solve this problem?

b. What is the balance outstanding after two years? How much principal and how much interest have been paid?

c. After two years, the interest rate falls to 6.75% p.a. What prepayment penalty would make it unattractive to prepay the loan?
Answer #1

a.

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 300,000, FV = 0, N = 120, I = 0.0725/12]

PMT = $3,522.03

Monthly Payment = $3,522.03

b.

Calculating Loan Balance after 2 years,

Using TVM Calculation,

FV = [PV = 300,000, N = 24, I = 0.0725/12, PMT = -3,522.03]

FV = $255,990.23

Loan Balance = $255,990.23

Principal Paid = 300,000 - 255,990.23 = $44,009.77

Interest Paid = 24(3,522.03) - 44,009.77 = $40,518.95